# NoisyNN: Exploring the Impact of Information Entropy Change in Learning Systems

### v3

@@ Appendix A Theoretical Foundations of Task Entropy
$\begin{split}H(\mathcal{T};\boldsymbol{Z})=&-\sum_{\boldsymbol{Y}\in\mathcal{Y}}p(\boldsymbol{Y}|\boldsymbol{Z})\log{p(\boldsymbol{Y}|\boldsymbol{Z})}\\ =&-\mathbb{E}[\log p(\boldsymbol{Y}|\boldsymbol{Z})]\\ =&-\mathbb{E}[\log[(2\pi)^{-k/2}|\boldsymbol{\Sigma_{Z}}|^{-1/2}\exp(-\frac{1}{2}\boldsymbol{(Y|Z-\mu_{Y|Z})}^{T}\boldsymbol{\Sigma}^{-1}_{\boldsymbol{Y|Z}}(\boldsymbol{Y|Z}-\boldsymbol{\mu_{Y|Z}}))]]\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|\end{split}$ (38)
@@
$\begin{split}\mathbb{E}[{(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})]=&\mathbb{E}[tr((\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})^{T}\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}))]\\ =&\mathbb{E}[tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})]\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}}){(\boldsymbol{Y|Z}-\boldsymbol{\mu}_{\boldsymbol{Y|Z}})}^{T})\\ =&tr(\Sigma^{-1}_{\boldsymbol{\boldsymbol{Y|Z}}}\Sigma_{\boldsymbol{\boldsymbol{Y|Z}}})\\ =&tr(\boldsymbol{I}_{k})\\ =&k\end{split}$ (39)
Therefore, for a specific set of embeddings, we can find that the task entropy is only related to the variance of the $\boldsymbol{Y|Z}$.
@@
$\left\{\begin{matrix}H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\boldsymbol{Z}\boldsymbol{\epsilon})\coloneqq H(\boldsymbol{Y};\boldsymbol{Z}\boldsymbol{\epsilon})-H(\boldsymbol{Z})&\boldsymbol{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (42)
@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{matrix}H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})=H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z})\end{matrix}.$ (43)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})=&H(\mathcal{T};\boldsymbol{Z})-H(\mathcal{T};\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Z})-(H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})-H(\boldsymbol{Z}))\\ =&H(\boldsymbol{Y};\boldsymbol{Z})-H(\boldsymbol{Y};\boldsymbol{Z}+\boldsymbol{\epsilon})\\ =&H(\boldsymbol{Y}|\boldsymbol{Z})+H(\boldsymbol{Z})-(H(\boldsymbol{Y}|\boldsymbol{Z+\boldsymbol{\epsilon}})+H(\boldsymbol{Z+\boldsymbol{\epsilon}}))\\ =&\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z}}|+\frac{1}{2}\log|\boldsymbol{\Sigma_{Z}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Y|Z+\epsilon}}|-\frac{1}{2}\log|\boldsymbol{\Sigma_{Z+\boldsymbol{\epsilon}}}|\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z}\Sigma_{ZY}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y}-\Sigma_{YZ}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{ZY}}|}\end{split}$ (44)
where $\boldsymbol{\Sigma_{Y|Z+\epsilon}}=\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}$. Since the Gaussian noise is independent of $\boldsymbol{Z}$ and $\boldsymbol{Y}$, we have $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=\boldsymbol{\Sigma_{YZ}}$. The corresponding proof is:
$\begin{split}\boldsymbol{\Sigma_{(Z+\epsilon)Y}}=&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}-\mu_{\boldsymbol{Z+\epsilon}}\right]\mathbb{E}\left[\boldsymbol{Y}-\mu_{\boldsymbol{Y}}\right]\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]-\mu_{\boldsymbol{Z+\epsilon}}\mathbb{E}\left[\boldsymbol{Y}\right]+\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z+\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mathbb{E}\left[\boldsymbol{(Z+\epsilon)}\right]\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]+\mathbb{E}\left[\boldsymbol{\epsilon}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{\epsilon}}\\ =&\mathbb{E}\left[\boldsymbol{Z}\boldsymbol{Y}\right]-\mu_{\boldsymbol{Y}}\mu_{\boldsymbol{Z}}\\ =&\boldsymbol{\Sigma_{ZY}}\end{split}$ (45)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})>0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\boldsymbol{\epsilon})\leq 0&if\ \frac{|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|}{|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|}\leq 1\end{matrix}\right.$ (46)
To find the relationship between $|\boldsymbol{\Sigma_{Z}}||\boldsymbol{\Sigma_{Y|Z}}|$ and $|\boldsymbol{\Sigma_{Z+\epsilon}}||\boldsymbol{\Sigma_{Y|Z+\epsilon}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the embeddings of the images are independent of each other, and so are the labels.
$\begin{split}\boldsymbol{\Sigma_{Y}}=&\mathbb{E}[(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})(\boldsymbol{Y}-\boldsymbol{{\mu}_{Y}})^{T}]\\ =&\mathbb{E}[\boldsymbol{Y}\boldsymbol{Y}^{T}]-\boldsymbol{{\mu}_{Y}}\boldsymbol{{\mu}_{Y}}^{T}\\ =&\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})\end{split}$ (47)
@@
The same procedure can be applied to $\boldsymbol{\Sigma_{Y(Z+\epsilon)}}$ and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$. Therefore, We can obtain that $\boldsymbol{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$,
$\boldsymbol{\Sigma_{Y(Z+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},Z_{1}+\epsilon),...,\mathrm{cov}(Y_{k},Z_{k}+\epsilon))$ (49)
and $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ is:
@@
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{Z_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{Z_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}+\sigma^{2}_{\epsilon}\boldsymbol{1}_{k}\end{split}$ (50)
@@
where $\boldsymbol{I}_{k}$ is a $k\times k$ identity matrix and $\boldsymbol{1}_{k}$ is a all ones $k\times k$ matrix. We use $\boldsymbol{U}$ to represent $\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\boldsymbol{I}_{k}$, and $\boldsymbol{u}$ to represent a all ones vector $[1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula [51] and Woodbury Formula [64], we can obtain the inverse of $\boldsymbol{{\Sigma}_{Z+\epsilon}}$ as:
@@
$\begin{split}\boldsymbol{{\Sigma}_{Z+\epsilon}^{-1}}=&(\boldsymbol{U}+\sigma^{2}_{\epsilon}\boldsymbol{uu}^{T})^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\boldsymbol{u}}\boldsymbol{U}^{-1}\boldsymbol{u}\boldsymbol{u}^{T}\boldsymbol{U}^{-1}\\ =&\boldsymbol{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}\boldsymbol{U}^{-1}\boldsymbol{1}_{k}\boldsymbol{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{Z_{1}}}-\frac{1}{\sigma^{4}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{2}}}-\frac{1}{\sigma^{4}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{Z_{k-1}}}-\frac{1}{\sigma^{4}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{k}}}-\frac{1}{\sigma^{4}_{Z_{k}}}\\ \end{bmatrix}\end{split}$ (51)
where $\boldsymbol{U}^{-1}=\mathrm{diag}((\sigma^{2}_{Z_{1}})^{-1},...,(\sigma^{2}_{Z_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$.
Therefore, substitute Equation 51 into ${|\boldsymbol{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$, we can obtain:

### v4

@@ Appendix A Theoretical Foundations of Task Entropy
$\begin{split}H(\mathcal{T};\bm{Z})=&-\sum_{\bm{Y}\in\mathcal{Y}}p(\bm{Y}|\bm{Z})\log{p(\bm{Y}|\bm{Z})}\\ =&-\mathbb{E}[\log p(\bm{Y}|\bm{Z})]\\ =&-\mathbb{E}\bigg{[}\log\big{[}(2\pi)^{-k/2}|\bm{\Sigma_{Z}}|^{-1/2}\\ &\quad\times\exp\big{(}-\frac{1}{2}(\bm{Y|Z}-\bm{\mu_{Y|Z}})^{T}\bm{\Sigma_{Y|Z}}^{-1}(\bm{Y|Z}-\bm{\mu_{Y|Z}})\big{)}\big{]}\bigg{]}\\ =&\frac{k}{2}\log(2\pi)+\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|\\ &+\frac{1}{2}\mathbb{E}[(\bm{Y|Z}-\bm{\mu}_{\bm{Y|Z}})^{T}\bm{\Sigma_{Y|Z}}^{-1}(\bm{Y|Z}-\bm{\mu}_{\bm{Y|Z}})]\\ =&\frac{k}{2}(1+\log(2\pi))+\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|\end{split}$ (29)
@@
$\begin{split}\mathbb{E}\left[(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})\right]=&\mathbb{E}\left[\mathrm{tr}\left((\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})\right)\right]\\ =&\mathbb{E}\left[\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\right)\right]\\ =&\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}\mathbb{E}\left[(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})(\bm{Y\mid Z}-\bm{\mu}_{\bm{Y\mid Z}})^{T}\right]\right)\\ =&\mathrm{tr}\left(\Sigma^{-1}_{\bm{Y\mid Z}}\Sigma_{\bm{Y\mid Z}}\right)\\ =&\mathrm{tr}(\bm{I}_{k})\\ =&k\end{split}$ (30)
Therefore, for a specific set of embeddings, we can find that the task entropy is only related to the variance of the $\bm{Y|Z}$.
@@
$\left\{\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ additive\ noise}\\ H(\mathcal{T};\bm{Z}\bm{\epsilon})\coloneqq H(\bm{Y};\bm{Z}\bm{\epsilon})-H(\bm{Z})&\bm{\epsilon}\ \mathrm{is\ multiplicative\ noise}\end{matrix}\right.$ (31)
@@ B.1 Inject Gaussian Noise in Latent Space
$\begin{matrix}H(\mathcal{T};\bm{Z}+\bm{\epsilon})=H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})\end{matrix}.$ (32)
@@
$\begin{split}\bigtriangleup S(\mathcal{T},\bm{\epsilon})=&\ H(\mathcal{T};\bm{Z})-H(\mathcal{T};\bm{Z}+\bm{\epsilon})\\ =&\ H(\bm{Y};\bm{Z})-H(\bm{Z})-\left(H(\bm{Y};\bm{Z}+\bm{\epsilon})-H(\bm{Z})\right)\\ =&H(\bm{Y};\bm{Z})-H(\bm{Y};\bm{Z}+\bm{\epsilon})\\ =&H(\bm{Y}|\bm{Z})+H(\bm{Z})-\left(H(\bm{Y}|\bm{Z}+\bm{\epsilon})+H(\bm{Z}+\bm{\epsilon})\right)\\ =&\frac{1}{2}\log|\bm{\Sigma_{Y|Z}}|+\frac{1}{2}\log|\bm{\Sigma_{Z}}|\\ &-\frac{1}{2}\log|\bm{\Sigma_{Y|Z+\epsilon}}|-\frac{1}{2}\log|\bm{\Sigma_{Z+\bm{\epsilon}}}|\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}\\ =&\frac{1}{2}\log\frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y}}-\Sigma_{YZ}\Sigma_{Z}^{-1}\Sigma_{ZY}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y}}-\Sigma_{YZ}\Sigma_{Z+\epsilon}^{-1}\Sigma_{ZY}|}\end{split}$ (33)
where $\bm{\Sigma_{Y|Z+\epsilon}}=\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}$. Since the Gaussian noise is independent of $\bm{Z}$ and $\bm{Y}$, we have $\bm{\Sigma_{Y(Z+\epsilon)}}=\bm{\Sigma_{(Z+\epsilon)Y}}=\bm{\Sigma_{YZ}}$. The corresponding proof is:
$\begin{split}\bm{\Sigma_{(Z+\epsilon)Y}}=&\mathbb{E}\left[\bm{(Z+\epsilon)}-\mu_{\bm{Z+\epsilon}}\right]\mathbb{E}\left[\bm{Y}-\mu_{\bm{Y}}\right]\\ =&\mathbb{E}\left[\bm{(Z+\epsilon)}\bm{Y}\right]-\mu_{\bm{Y}}\mathbb{E}\left[\bm{(Z+\epsilon)}\right]-\mu_{\bm{Z+\epsilon}}\mathbb{E}\left[\bm{Y}\right]+\mu_{\bm{Y}}\mu_{\bm{Z+\epsilon}}\\ =&\mathbb{E}\left[\bm{(Z+\epsilon)}\bm{Y}\right]-\mu_{\bm{Y}}\mathbb{E}\left[\bm{(Z+\epsilon)}\right]\\ =&\mathbb{E}\left[\bm{Z}\bm{Y}\right]+\mathbb{E}\left[\bm{\epsilon}\bm{Y}\right]-\mu_{\bm{Y}}\mu_{\bm{Z}}-\mu_{\bm{Y}}\mu_{\bm{\epsilon}}\\ =&\mathbb{E}\left[\bm{Z}\bm{Y}\right]-\mu_{\bm{Y}}\mu_{\bm{Z}}\\ =&\bm{\Sigma_{ZY}}\end{split}$ (34)
@@
$\left\{\begin{matrix}\bigtriangleup S(\mathcal{T},\bm{\epsilon})>0&if\ \frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}>1\\ \bigtriangleup S(\mathcal{T},\bm{\epsilon})\leq 0&if\ \frac{|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|}{|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|}\leq 1\end{matrix}\right.$ (35)
To find the relationship between $|\bm{\Sigma_{Z}}||\bm{\Sigma_{Y|Z}}|$ and $|\bm{\Sigma_{Z+\epsilon}}||\bm{\Sigma_{Y|Z+\epsilon}}|$, we need to determine the subterms in each of them. As we mentioned in the previous section, the embeddings of the images are independent of each other, and so are the labels.
$\begin{split}\bm{\Sigma_{Y}}=&\mathbb{E}[(\bm{Y}-\bm{{\mu}_{Y}})(\bm{Y}-\bm{{\mu}_{Y}})^{T}]\\ =&\mathbb{E}[\bm{Y}\bm{Y}^{T}]-\bm{{\mu}_{Y}}\bm{{\mu}_{Y}}^{T}\\ =&\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})\end{split}$ (36)
@@
The same procedure can be applied to $\bm{\Sigma_{Y(Z+\epsilon)}}$ and $\bm{{\Sigma}_{Z+\epsilon}}$. Therefore, We can obtain that $\bm{\Sigma_{Y}}=\mathrm{diag}(\sigma^{2}_{Y_{1}},...,\sigma^{2}_{Y_{k}})$,
$\bm{\Sigma_{Y(Z+\epsilon)}}=\mathrm{diag}(\mathrm{cov}(Y_{1},Z_{1}+\epsilon),...,\mathrm{cov}(Y_{k},Z_{k}+\epsilon))$ (38)
and $\bm{{\Sigma}_{Z+\epsilon}}$ is:
@@
$\begin{split}\bm{{\Sigma}_{Z+\epsilon}}=&\begin{bmatrix}\sigma^{2}_{Z_{1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{2}}+\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \vdots&\vdots&&\vdots&\vdots\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{Z_{k-1}}+\sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}\\ \sigma^{2}_{\epsilon}&\sigma^{2}_{\epsilon}&...&\sigma^{2}_{\epsilon}&\sigma^{2}_{Z_{k}}+\sigma^{2}_{\epsilon}\\ \end{bmatrix}\\ =&\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\bm{I}_{k}+\sigma^{2}_{\epsilon}\bm{1}_{k}\end{split}$ (39)
@@
where $\bm{I}_{k}$ is a $k\times k$ identity matrix and $\bm{1}_{k}$ is a all ones $k\times k$ matrix. We use $\bm{U}$ to represent $\mathrm{diag}(\sigma^{2}_{Z_{1}},...,\sigma^{2}_{Z_{k}})\bm{I}_{k}$, and $\bm{u}$ to represent a all ones vector $[1,...,1]^{T}$. Thanks to the Sherman–Morrison Formula Sherman & Morrison (1949) and Woodbury Formula Woodbury (1950), we can obtain the inverse of $\bm{{\Sigma}_{Z+\epsilon}}$ as:
@@
$\begin{split}\bm{{\Sigma}_{Z+\epsilon}^{-1}}=&(\bm{U}+\sigma^{2}_{\epsilon}\bm{uu}^{T})^{-1}\\ =&\bm{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sigma^{2}_{\epsilon}\bm{u}^{T}\bm{U}^{-1}\bm{u}}\bm{U}^{-1}\bm{u}\bm{u}^{T}\bm{U}^{-1}\\ =&\bm{U}^{-1}-\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}\bm{U}^{-1}\bm{1}_{k}\bm{U}^{-1}\\ =&\lambda\begin{bmatrix}\frac{1}{\lambda\sigma^{2}_{Z_{1}}}-\frac{1}{\sigma^{4}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{2}}}-\frac{1}{\sigma^{4}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{2}}\sigma^{2}_{Z_{k}}}\\ \vdots&\vdots&&\vdots&\vdots\\ -\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{2}}}&...&\frac{1}{\lambda\sigma^{2}_{Z_{k-1}}}-\frac{1}{\sigma^{4}_{Z_{k-1}}}&-\frac{1}{\sigma^{2}_{Z_{k-1}}\sigma^{2}_{Z_{k}}}\\ -\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{1}}}&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{2}}}&...&-\frac{1}{\sigma^{2}_{Z_{k}}\sigma^{2}_{Z_{k-1}}}&\frac{1}{\lambda\sigma^{2}_{Z_{k}}}-\frac{1}{\sigma^{4}_{Z_{k}}}\\ \end{bmatrix}\end{split}$ (40)
where $\bm{U}^{-1}=\mathrm{diag}((\sigma^{2}_{Z_{1}})^{-1},...,(\sigma^{2}_{Z_{k}})^{-1})$ and $\lambda=\frac{\sigma^{2}_{\epsilon}}{1+\sum_{i=1}^{k}\frac{1}{\sigma^{2}_{Z_{i}}}}$.
Therefore, substitute Equation 40 into ${|\bm{\Sigma_{Y}-\Sigma_{Y(Z+\epsilon)}{\Sigma}^{-1}_{Z+\epsilon}\Sigma_{(Z+\epsilon)Y}}|}$, we can obtain: